### v1

@@ Abstract
Understanding the nature of primordial fluctuations is critical to our comprehension of the Universe’s early stages. While these fluctuations are known to be nearly scale-invariant, quasi-adiabatic, and nearly Gaussian on large scales, their behavior at smaller scales remains less well-defined and may offer insights into new physics. Recent observations by the NANOGrav, PPTA, EPTA, and CPTA collaborations suggest the presence of a stochastic gravitational wave background, which, while consistent with the contribution from supermassive black hole binaries, also opens the possibility of probing new physics. This paper explores whether this signal could stem from primordial isocurvature and adiabatic fluctuations. We adopt parameterized spectra for both types of fluctuations to fit the observations from the latest NANOGrav data. Furthermore, we employ Bayesian analysis to assess the distinguishability of these models in light of current PTA sensitivities. Our findings indicate that with the capabilities, PTAs cannot conclusively differentiate between isocurvature and adiabatic fluctuations.
@@ 1 Introduction
Recently, several PTA collaborations, including the North American Nanohertz Observatory for Gravitational Waves (NANOGrav) [16, 17], the Parkers PTA (PPTA) [18, 19], the European PTA (EPTA) in conjunction with the Indian PTA (InPTA) [20, 21], and the Chinese PTA (CPTA) [22], have collectively unveiled compelling evidence supporting the existence of a stochastic gravitational wave background (SGWB) within the frequency range of approximately $1-10\,\mathrm{nHz}$, which attracts lots of implications [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61]. While the inferred amplitude and spectrum of the GW background align with astrophysical expectations for a signal originating from the population of supermassive black hole binaries (SMBHBs), the possibility of new physics within this observational window remains a tantalizing prospect. One plausible explanation for the observed SGWB is that it arises from IGWs stemming from primordial fluctuations [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99]. Furthermore, cosmological phenomena such as first-order phase transitions [100, 101, 102, 103, 104, 105, 106, 107, 108] and the presence of topological defects like cosmic strings [109, 110, 111, 112, 113, 114, 115, 116] and domain walls [117, 118, 119] could also serve as sources for generating the SGWB. The SGWB generated by these cosmological sources holds immense potential for advancing our understanding of physics beyond the Standard Model and providing invaluable insights into the early stages of the Universe.
In this study, our goal is to investigate the potential explanations for the observed signals through IGWs stemming from both primordial isocurvature and adiabatic fluctuations. To achieve this, we utilize parameterized primordial spectra for both types of fluctuations to effectively match the latest observations reported by the NANOGrav. Moreover, we calculate the Bayes factors to compare different models. Our analysis indicates that, with the current sensitivity of PTAs, it is challenging to differentiate between the contributions of isocurvature and adiabatic fluctuations. The structure of this paper is outlined as follows: In Section 2, we introduce the general formalisms for IGWs arising from both types of primordial fluctuations. Section 3 details the methodology employed in our data analysis and presents the findings using the latest NANOGrav data set, along with the computed Bayes factors for comparing different models. We conclude by summarizing our results and discussing their implications in Section 4.
@@ 2 IGWs from primordial isocurvature fluctuations and primordial adiabatic fluctuations
After generation, the energy density of GWs evolves in the same way as radiation. Utilizing this property, it is straightforward to determine the energy density of GWs at present. Taking the late-time limit during the radiation-dominated era, the relation of the GW energy density at present and the generation is given by [126]
@@
$\mathcal{T}_{s}(\kappa,u,v)=\frac{1}{3}\left(\frac{4v^{2}-(1-u^{2}+v^{2})^{2}}{4uv}\right)^{2}\left(I_{c,\infty}^{2}(\kappa,u,v)+I_{s,\infty}^{2}(\kappa,u,v)\right)\,,$ (6)
@@
To maintain a high level of model independence, we adopt an approach that avoids selecting a specific inflation model capable of generating an enhanced spectrum $\mathcal{P}_{\zeta/s}$. Instead, we focus on three characteristic templates for $\mathcal{P}_{\zeta/s}$ that encompass a range of possibilities typically observed in realistic models. By disregarding the microphysics of inflation, we aim to provide a broader analysis that accommodates various scenarios. The following templates are considered:
@@ 3 Data analyses and results
Table 1: Parameters along with their respective prior distributions utilized in the Bayesian inference analysis, where $\mathcal{U}$ denotes a uniform distribution. Results are reported as the median and the $90\%$ credible interval, with symmetrical tails.
Model Parameter Description Prior Results
$M_{1}$ adiabatic delta
$\log_{10}A$ Amplitude of the power spectrum. ${\mathcal{U}}(-5,5)$ $0.56^{+1.83}_{-1.99}$
$\log_{10}(f_{*}/\mathrm{Hz})$ Pivot frequency. ${\mathcal{U}}(-15,-2)$ $-4.28^{+2.02}_{-2.45}$
$M_{2}$ adiabatic lognormal
$\log_{10}A$ Amplitude of the power spectrum. ${\mathcal{U}}(-2,2)$ $0.21^{+1.47}_{-1.52}$
$\log_{10}(f_{*}/\mathrm{Hz})$ Pivot frequency. ${\mathcal{U}}(-8,-2)$ $-4.97^{+2.58}_{-2.31}$
$\Delta$ Width of the power spectrum. ${\mathcal{U}}(0.01,5)$ $2.35^{+1.70}_{-1.98}$
$M_{3}$ adiabatic box
$\log_{10}A$ Amplitude of the power spectrum. ${\mathcal{U}}(-3,3)$ $-1.26^{+2.28}_{-0.61}$
$\log_{10}(f_{\mathrm{min}}/\mathrm{Hz})$ Minimum frequency of the power spectrum. ${\mathcal{U}}(-9,-4)$ $-7.37^{+1.86}_{0.04}$
$\log_{10}(f_{\mathrm{max}}/\mathrm{Hz})$ Maximum frequency of the power spectrum. ${\mathcal{U}}(-9,-2)$ $-4.60^{+2.34}_{-3.45}$
$M_{4}$ isocurvature delta
$\log_{10}A$ Amplitude of the power spectrum. ${\mathcal{U}}(1,50)$ $25.3^{+5.9}_{-6.9}$
$\log_{10}(f_{*}/\mathrm{Hz})$ Pivot frequency. ${\mathcal{U}}(-15,-2)$ $-4.25^{+2.03}_{-2.44}$
$M_{5}$ isocurvature lognormal
$\log_{10}A$ Amplitude of the power spectrum. ${\mathcal{U}}(10,50)$ $21.7^{+5.2}_{-4.3}$
$\log_{10}(f_{*}/\mathrm{Hz})$ Pivot frequency. ${\mathcal{U}}(-8,-2)$ $-4.95^{+2.64}_{-2.18}$
$\Delta$ Width of the power spectrum. ${\mathcal{U}}(0.01,5)$ $1.16^{+0.79}_{-1.11}$
$M_{6}$ isocurvature box
$\log_{10}A$ Amplitude of the power spectrum. ${\mathcal{U}}(10,30)$ $18.3^{+4.9}_{-1.3}$
$\log_{10}(f_{\mathrm{min}}/\mathrm{Hz})$ Pivot frequency. ${\mathcal{U}}(-9,-4)$ $-7.10^{+1.54}_{-0.45}$
$\log_{10}(f_{\mathrm{max}}/\mathrm{Hz})$ Width of the power spectrum. ${\mathcal{U}}(-9,-2)$ $-4.41^{+2.17}_{-2.44}$
@@
$f=\frac{k}{2\pi}\simeq 1.6\,\mathrm{nHz}\left(\frac{k}{10^{6}\,\mathrm{Mpc}^{-1}}\right).$ (16)
which provides the necessary link for interpreting spectral features in terms of their spatial scales. In this work, we consider six models: the adiabatic delta model ($M_{1}$), adiabatic lognormal model ($M_{2}$), adiabatic box model ($M_{3}$), isocurvature delta model ($M_{4}$), isocurvature lognormal model ($M_{5}$), and isocurvature box model ($M_{6}$). The selected priors for the model parameters are comprehensively listed in Table 1.
Figure 1: Marginalized one- and two-dimensional posterior distributions for the model parameters of the adiabatic delta (upper panel) and isocurvature delta (lower panel) models, derived from the NANOGrav 15-year data set. The contours in the 2D distributions represent the $1\sigma$, $2\sigma$, and $3\sigma$ credible intervals, respectively.
The posterior distributions for the parameters of the models $M_{1}$ through $M_{6}$ are illustrated in Figs. 1-3. Our analysis indicates that a successful explanation of the NANOGrav signal with IGWs requires a substantial amplitude $A$ for the primordial scalar power spectrum. This inference is supported by computing the lower limits of the $95\%$ Bayesian credible intervals for the amplitude $A$, derived from the one-dimensional marginalized posteriors over the regions of the highest posterior density. For models $M_{1}$ through $M_{6}$, we find the following lower limits: $\log_{10}A\gtrsim-1.43$, $\log_{10}A\gtrsim-1.31$, $\log_{10}A\gtrsim-1.88$, $\log_{10}A\gtrsim 18.3$, $\log_{10}A\gtrsim 17.4$, and $\log_{10}A\gtrsim 17.0$, respectively. Additionally, the amplification in scalar perturbations must be confined to specific scales to ensure that the resulting IGW signal is within the sensitivity range of PTAs. This constraint imposes specific bounds on the characteristic frequencies $f_{\mathrm{min}}$, $f_{\mathrm{max}}$, and $f_{*}$, which are evident from Figs. 1-3 and are concisely presented in Table 1.
@@
This behavior is further corroborated by the flat trends observed in the 2D posterior distributions of the amplitude $A$ with both $f_{\mathrm{min}}$ and $f_{*}$, as illustrated in Figs. 1-3. These flat directions in the posterior space indicate that increasing $A$ simultaneously with $f_{\mathrm{min}}$ or $f_{*}$ shifts the GW spectrum’s peak to higher frequencies. Crucially, this shift does not alter the low-frequency tail within the PTA band, thereby preserving the fit’s consistency with the NANOGrav data.
@@
The calculation of Bayes factors among six competing models yields the following results:
$\mathrm{BF}_{ij}=\left(\begin{array}[]{ccccccc}&M_{1}&M_{2}&M_{3}&M_{4}&M_{5}&M_{6}\\ M_{1}&1.00&0.54&0.87&1.78&8.50&2.76\\ M_{2}&1.86&1.00&1.62&3.32&15.85&5.15\\ M_{3}&1.18&0.62&1.00&2.04&9.76&3.17\\ M_{4}&0.56&0.30&0.49&1.00&4.77&1.55\\ M_{5}&0.12&0.06&0.10&0.21&1.00&0.32\\ M_{6}&0.36&0.19&0.32&0.64&3.07&1.00\end{array}\right).$ (19)
Upon analysis, model $M_{2}$ exhibits the highest Bayes factor, in contrast to model $M_{5}$, which has the lowest. Remarkably, the Bayes factor comparing $M_{2}$ to $M_{5}$ is $\mathrm{BF}_{25}=15.85$, indicating positive evidence rather than significant evidence in favor of $M_{2}$ over $M_{5}$. Therefore, the interpretation of the NANOGrav 15-year dataset concerning the common-spectrum process and the associated Hellings-Downs correlation remains nuanced. The Bayes factors, when considered collectively, provide evidence that is not strong enough to decisively favor any single model over the others. Consequently, selecting the optimal model among $M_{1}$ through $M_{6}$ is challenging, given the narrow range of their Bayes factors and the level of evidence they present.

### v2

@@ Abstract
Understanding the nature of primordial fluctuations is pivotal to unraveling the Universe’s early evolution. While these fluctuations are observed to be nearly scale-invariant, quasi-adiabatic, and Gaussian on large scales, their small-scale behavior remains poorly constrained, offering a potential window into new physics. Recent detections of a stochastic gravitational wave background in the nanohertz frequency range by pulsar timing arrays (PTAs), including NANOGrav, PPTA, EPTA+InPTA, and CPTA, align with astrophysical predictions from supermassive black hole binaries but could also encode signatures of primordial phenomena. We investigate whether the observed signal originates from primordial isocurvature or adiabatic fluctuations by fitting them to the latest NANOGrav dataset. Through comprehensive Bayesian model comparison, we evaluate the distinguishability of these scenarios given current PTA sensitivities. Our results demonstrate that existing data cannot conclusively differentiate between isocurvature and adiabatic sources, highlighting the need for enhanced observational capabilities to probe the primordial universe at small scales.
@@ 1 Introduction
Recently, several PTA collaborations, including the North American Nanohertz Observatory for Gravitational Waves (NANOGrav) [16, 17], the Parkers PTA (PPTA) [18, 19], the European PTA (EPTA) in conjunction with the Indian PTA (InPTA) [20, 21], and the Chinese PTA (CPTA) [22], have collectively unveiled compelling evidence supporting the existence of a stochastic gravitational wave background (SGWB) within the frequency range of approximately $1-10\,\mathrm{nHz}$, which attracts lots of implications [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61]. While the inferred amplitude and spectrum of the SGWB align with astrophysical expectations for a signal originating from the population of supermassive black hole binaries (SMBHBs), the possibility of new physics within this observational window remains a tantalizing prospect. One plausible explanation for the observed SGWB is that it arises from IGWs stemming from primordial fluctuations [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99]. Furthermore, cosmological phenomena such as first-order phase transitions [100, 101, 102, 103, 104, 105, 106, 107, 108] and the presence of topological defects like cosmic strings [109, 110, 111, 112, 113, 114, 115, 116] and domain walls [117, 118, 119] could also serve as sources for generating the SGWB. The SGWB generated by these cosmological sources holds immense potential for advancing our understanding of physics beyond the Standard Model and providing invaluable insights into the early stages of the Universe.
In this study, our goal is to investigate the potential explanations for the observed signals through IGWs stemming from both primordial isocurvature and adiabatic fluctuations. To achieve this, we utilize parameterized primordial spectra for both types of fluctuations to effectively match the latest observations reported by the NANOGrav. Moreover, we calculate the Bayes factors to compare different models. Our analysis indicates that, with the current sensitivity of PTAs, it is challenging to differentiate between the contributions of isocurvature and adiabatic fluctuations. The structure of this paper is outlined as follows: In Section 2, we introduce the general formalisms for IGWs arising from both types of primordial fluctuations. In Section 3, we provide an overview of the abundance of PBH from both types of primordial fluctuations. Section 4 details the methodology employed in our data analysis and presents the findings using the latest NANOGrav data set, along with the computed Bayes factors for comparing different models. We conclude by summarizing our results and discussing their implications in Section 5.
@@ 2 IGWs from primordial isocurvature fluctuations and primordial adiabatic fluctuations
After generation, the energy density of GWs evolves in the same way as radiation. Utilizing this property, it is straightforward to determine the energy density of GWs at present. Taking the late-time limit during the radiation-dominated era, the relation of the GW energy density at present ($\Omega_{\rm GW}$) and the generation ($\Omega_{\rm GW,c}$) is given by [126]
@@
$\mathcal{T}_{s}(\kappa,u,v)=\frac{1}{3}\left(\frac{4v^{2}-(1-u^{2}+v^{2})^{2}}{4uv}\right)^{2}\left(I_{c,\infty}^{2}+I_{s,\infty}^{2}\right)\,,$ (6)
@@
3 PBHs from primordial isocurvature fluctuations and primordial adiabatic fluctuations
The production of IGWs depends on amplified scalar perturbations at small scales, which can concurrently result in the formation of PBHs. This section introduces the formalisms for PBHs originating from sizable primordial isocurvature and adiabatic fluctuations. The total fraction of PBHs in the present-day dark matter can be defined as
$f_{\mathrm{PBH}}\equiv\frac{\Omega_{\mathrm{PBH}}}{\Omega_{\mathrm{DM}}}=\int\frac{dM}{M}f(M),$ (9)
where $\Omega_{\mathrm{DM}}$ represents the dark matter fraction and $f(M)$ denotes the PBH mass function. The horizon mass $M_{H}$ and the wavenumber $k$ are related by the following expression:
$M_{H}\simeq\gamma M_{\odot}\left(\frac{g_{*,s}(T)}{10.75}\right)^{-2/3}\left(\frac{g_{*,r}(T)}{10.75}\right)^{1/2}\left(\frac{4.2\times 10^{6}\mathrm{Mpc}^{-1}}{k}\right)^{2}.$ (10)
Table 1: Parameters along with their respective prior distributions utilized in the Bayesian inference analysis, where $\mathcal{U}$ denotes a uniform distribution. Additional physical constraints are imposed requiring $f_{\rm PBH}\leqslant 1$ to ensure PBH abundance does not exceed dark matter density (see Sec. 3 for more details), as well as that the constraints from CMB [128] and BBN [129]. Results are reported as the median and the $90\%$ credible interval, with symmetrical tails.
Model Parameter Description Prior Results
$M_{0}$ SMBHB
$\log_{10}A_{\mathrm{SMBHB}}$ Amplitude of the power spectrum. ${\mathcal{U}}(-15,-12)$ $-14.58^{+0.09}_{-0.15}$
$M_{1}$ Adiabatic Delta
$\log_{10}A$ Amplitude of the power spectrum. ${\mathcal{U}}(-3,0)$ $-1.46^{+0.21}_{-0.34}$
$\log_{10}(f_{*}/\mathrm{Hz})$ Pivot frequency. ${\mathcal{U}}(-8,-5)$ $-6.74^{+0.32}_{-0.51}$
$M_{2}$ Adiabatic Lognormal
$\log_{10}A$ Amplitude of the power spectrum. ${\mathcal{U}}(-3,0)$ $-0.81^{+0.48}_{-0.70}$
$\log_{10}(f_{*}/\mathrm{Hz})$ Pivot frequency. ${\mathcal{U}}(-8,-2)$ $-6.71^{+1.11}_{-0.79}$
$\Delta$ Width of the power spectrum. ${\mathcal{U}}(0.01,5)$ $1.18^{+1.84}_{-1.02}$
$M_{3}$ Adiabatic Box
$\log_{10}A$ Amplitude of the power spectrum. ${\mathcal{U}}(-3,0)$ $-1.46^{+0.77}_{-0.47}$
$\log_{10}(f_{\mathrm{min}}/\mathrm{Hz})$ Minimum frequency of the power spectrum. ${\mathcal{U}}(-9,-4)$ $-7.57^{+0.67}_{-0.59}$
$\log_{10}(f_{\mathrm{max}}/\mathrm{Hz})$ Maximum frequency of the power spectrum. ${\mathcal{U}}(-8,-2)$ $-5.45^{+3.02}_{-1.60}$
$M_{4}$ Isocurvature Delta
$\log_{10}A$ Amplitude of the power spectrum. ${\mathcal{U}}(16,20)$ $17.9^{+0.7}_{-1.2}$
$\log_{10}(f_{*}/\mathrm{Hz})$ Pivot frequency. ${\mathcal{U}}(-8,-6)$ $-6.88^{+0.30}_{-0.46}$
$M_{5}$ Isocurvature Lognormal
$\log_{10}A$ Amplitude of the power spectrum. ${\mathcal{U}}(10,30)$ $21.1^{+4.4}_{-3.9}$
$\log_{10}(f_{*}/\mathrm{Hz})$ Pivot frequency. ${\mathcal{U}}(-8,-2)$ $-4.88^{+2.63}_{-2.37}$
$\Delta$ Width of the power spectrum. ${\mathcal{U}}(0.01,5)$ $1.39^{+0.64}_{-1.18}$
$M_{6}$ Isocurvature box
$\log_{10}A$ Amplitude of the power spectrum. ${\mathcal{U}}(15,25)$ $18.0^{+2.6}_{-1.1}$
$\log_{10}(f_{\mathrm{min}}/\mathrm{Hz})$ Pivot frequency. ${\mathcal{U}}(-8,-6)$ $-7.18^{+0.83}_{-0.37}$
$\log_{10}(f_{\mathrm{max}}/\mathrm{Hz})$ Width of the power spectrum. ${\mathcal{U}}(-7,-2)$ $-4.54^{+2.27}_{-2.13}$
According to Refs. [130, 131], the mass function of PBHs arising from primordial adiabatic fluctuations can be expressed as
$f(M)\simeq\gamma^{3/2}\left(\frac{\beta(M)}{1.6\times 10^{-9}}\right)\left(\frac{10.75}{g_{*,r}(T)}\right)^{-3/4}\left(\frac{10.75}{g_{*,s}(T)}\right)\left(\frac{0.12}{\Omega_{\mathrm{DM}}h^{2}}\right)\left(\frac{M_{\odot}}{M}\right)^{1/2},$ (11)
where $\gamma\simeq 0.4$ is the ratio between the PBH mass $M$ and the horizon mass $M_{H}$ [132], i.e., $M=\gamma M_{H}$, and $\beta(M)$ is the the PBH production rate. Assuming a Gaussian distribution for the density perturbations, the PBH production rate can be calculated by integrating the Gaussian distribution of perturbations, namely
$\beta(M)=\int_{\delta_{c}}\frac{d\delta}{\sqrt{2\pi}\sigma_{\zeta}(M)}\exp\left(-\frac{\delta^{2}}{2\sigma_{\zeta}^{2}(M)}\right),$ (12)
where $\delta_{c}\simeq 0.45$ is the critical density contrast equired for PBH formation [133], and $\sigma_{\zeta}^{2}(M)$ is the variance of the density contrast, which in the radiation domination era can be given by
$\sigma_{\zeta}^{2}(M)=\frac{16}{81}\int\frac{dq}{q}\left(\frac{q}{k(M)}\right)^{4}W^{2}\left(\frac{q}{k(M)}\right)T^{2}\left(q,k^{-1}(M)\right)\mathcal{P}_{\zeta}(q).$ (13)
Here $W(x)=e^{-x^{2}/2}$ is the Fourier transform of a volume-normalized Gaussian window smoothing function, and
$\mathcal{T}\left(q,k^{-1}\right)=3\left[\sin\left(\frac{q}{\sqrt{3}k}\right)-\left(\frac{q}{\sqrt{3}k}\right)\cos\left(\frac{q}{\sqrt{3}k}\right)\right]\left(\frac{q}{\sqrt{3}k}\right)^{-3}$ (14)
denotes the transfer function. Following Refs. [134, 125], the mass function of PBHs originating from primordial isocurvature fluctuations can be expressed as
$f(M)\simeq\frac{3^{q/2}k_{\text{eq}}^{2q}\sigma_{s}^{2q+1}(M)b}{2^{q}k^{2q}\xi},$ (15)
where $b\simeq 0.02$, $q\simeq 13/2$ , $\xi\simeq 1$ are constants which encode the PBH formation probability [134], and the variance of the fluctuation $\sigma_{s}(M)$ is given by
$\sigma_{s}(M)\simeq\int d\ln(q)\left(\frac{q}{k(M)}\right)^{4}W^{2}(\frac{q}{k(M)})\mathcal{P}_{s}(q).$ (16)
@@ 3 PBHs from primordial isocurvature fluctuations and primordial adiabatic fluctuations
To maintain a high level of model independence, we adopt an approach that avoids selecting a specific inflation model capable of generating an enhanced spectrum $\mathcal{P}_{\zeta/s}$. Instead, we focus on three characteristic templates for $\mathcal{P}_{\zeta/s}$ that encompass a range of possibilities typically observed in realistic models. By disregarding the microphysics of inflation, we aim to provide a broader analysis that accommodates various scenarios. The following functional forms of power spectrum are considered:
@@
An illustration of the energy density spectrum of IGWs as a function of frequency $f$ for different scenarios can be found in Fig. 1.
Figure 1: Energy density spectrum of IGWs ($\Omega_{\rm GW}$) as a function of frequency $f$. The gray violin plots represent the free spectral measurements from the NANOGrav 15-year dataset. Colored lines show theoretical predictions for adiabatic (solid) and isocurvature (dashed) scenarios with model parameters fixed to their median posterior values.
@@ 4 Data analyses and results
$f=\frac{k}{2\pi}\simeq 1.6\,\mathrm{nHz}\left(\frac{k}{10^{6}\,\mathrm{Mpc}^{-1}}\right),$ (24)
which provides the necessary link for interpreting spectral features in terms of their spatial scales. In this work, we consider six models for IGWs: the adiabatic delta model ($M_{1}$), adiabatic lognormal model ($M_{2}$), adiabatic box model ($M_{3}$), isocurvature delta model ($M_{4}$), isocurvature lognormal model ($M_{5}$), and isocurvature box model ($M_{6}$). In addition, we consider the SMBHB model ($M_{0}$). The SGWB produced by SMBHBs can be approximated by a power-law spectrum [139]:
$\Omega_{\mathrm{GW}}(f)=\frac{2\pi^{2}A_{\mathrm{SMBHB}}^{2}}{3H_{0}^{2}}\left(\frac{f}{f_{\mathrm{yr}}}\right)^{5-\gamma}f_{\mathrm{yr}}^{2},$ (25)
where $A_{\rm SMBHB}$ is the amplitude of the characteristic strain measured at the reference frequency $f_{\rm yr}\equiv 1\,{\rm yr}^{-1}$, and $\gamma=13/3$ is the expected power-law spectral index for a population of circular, GW-driven binaries. The selected priors for the model parameters are comprehensively listed in Table 1. Moreover, physical constraints are imposed requiring $f_{\rm PBH}\leqslant 1$ to ensure PBH abundance does not exceed dark matter density, as well as that the constraints from CMB [128] and BBN [129].
Figure 2: Marginalized one- and two-dimensional posterior distributions for model parameters in the adiabatic delta (upper panel) and isocurvature delta (lower panel) scenarios, analyzed using the NANOGrav 15-year dataset. Blue contours show results for pure IGW models, while orange contours represent the combined IGW+SMBHB analysis. The contours in two-dimensional distributions denote the $68\%$, $95\%$, and $99.7\%$ credible regions ($1\sigma$, $2\sigma$, and $3\sigma$, respectively).
The posterior distributions for the parameters of the models $M_{1}$ through $M_{6}$ are illustrated in Figs. 2-4. Our analysis indicates that a successful explanation of the NANOGrav signal with IGWs requires a substantial amplitude $A$ for the primordial scalar power spectrum. This inference is supported by computing the lower limits of the $95\%$ Bayesian credible intervals for the amplitude $A$, derived from the one-dimensional marginalized posteriors over the regions of the highest posterior density. For models $M_{1}$ through $M_{6}$, we find the following lower limits: $\log_{10}A\gtrsim-1.80$, $\log_{10}A\gtrsim-1.51$, $\log_{10}A\gtrsim-1.93$, $\log_{10}A\gtrsim 16.7$, $\log_{10}A\gtrsim 17.2$, and $\log_{10}A\gtrsim 16.9$, respectively. Additionally, the amplification in scalar perturbations must be confined to specific scales to ensure that the resulting IGW signal is within the sensitivity range of PTAs. Additionally, the parameter space is constrained by multiple physical requirements: the avoidance of PBH overproduction, consistency with CMB and BBN bounds. These constraint imposes specific bounds on the characteristic frequencies $f_{\mathrm{min}}$, $f_{\mathrm{max}}$, and $f_{*}$, which are evident from Figs. 2-4 and are concisely presented in Table 1.
@@
The calculation of Bayes factors among seven competing models yields the following results:
$\mathrm{BF}_{ij}=\left(\begin{array}[]{cccccccc}&{M_{0}}&{M_{1}}&{M_{2}}&{M_{3}}&{M_{4}}&{M_{5}}&{M_{6}}\\ {M_{0}}&{1.00}&{0.12}&{0.16}&{0.09}&{0.17}&{1.23}&{0.08}\\ {M_{1}}&{8.56}&{1.00}&{1.33}&{0.80}&{1.45}&{10.51}&{0.67}\\ {M_{2}}&{6.45}&{0.75}&{1.00}&{0.60}&{1.10}&{7.92}&{0.51}\\ {M_{3}}&{10.70}&{1.25}&{1.66}&{1.00}&{1.82}&{13.13}&{0.84}\\ {M_{4}}&{5.89}&{0.69}&{0.91}&{0.55}&{1.00}&{7.23}&{0.46}\\ {M_{5}}&{0.81}&{0.10}&{0.13}&{0.08}&{0.14}&{1.00}&{0.06}\\ {M_{6}}&{12.76}&{1.49}&{1.98}&{1.19}&{2.17}&{15.66}&{1.00}\end{array}\right).$ (28)
It can be seen that the IGW models show modest preference over the pure SMBHB model, with Bayes factors ranging from $0.81$ to $12.76$. Model $M_{6}$ exhibits the highest Bayes factor, while model $M_{5}$ yields the lowest. The Bayes factor comparing these models, $\mathrm{BF}_{65}=15.66$, indicates positive but not strong evidence favoring $M_{6}$ over $M_{5}$ on the Jeffreys scale. The interpretation of the NANOGrav 15-year dataset regarding the common-spectrum process and the associated Hellings-Downs correlation remains subtle, as the collective Bayes factors do not provide decisive evidence for any particular model. The relatively narrow range of Bayes factors among models $M_{0}$ through $M_{6}$ precludes a definitive model selection.
[…]
$M_{1}$ $0.14$
$M_{2}$ $0.24$
$M_{3}$ $0.15$
$M_{4}$ $0.17$
$M_{5}$ $0.53$
$M_{6}$ $0.54$
We further analyze the statistical evidence for an additional SMBHB component by computing Bayes factors between the combined IGW+SMBHB models and their corresponding pure IGW counterparts. The results are presented in Table 3. The obtained Bayes factors are consistently less than unity ($\mathrm{BF}<1$), indicating that the NANOGrav 15-year dataset does not support the presence of an SMBHB contribution beyond the IGW signal. Furthermore, the inclusion of SMBHB components results in increased posterior widths for IGW parameters, an effect expected from the additional degrees of freedom in the combined models.